OpenStudy (anonymous):

Isadora bought 4 trees to plant in her backyard and spent no less than eighty-four dollars altogether. How many dollars could each tree have cost? Let t represent how many dollars each tree cost. t is greater than eighty t is greater than or equal to eighty t is greater than twenty-one t is greater than or equal to twenty-one

OpenStudy (anonymous):

4 divide 4 is 1

OpenStudy (anonymous):

8 divided by 4?

OpenStudy (nincompoop):

\[4t \ge 84\]

OpenStudy (anonymous):

t is greater than or equal to eighty

OpenStudy (nincompoop):

the question is: how much could EACH tree cost? meaning you will have to divide 4t by 4 and also 84 by 4 this gives you: \[\frac{ 4t }{ 4 } \ge \frac{ 84 }{ 4 }\]

OpenStudy (anonymous):

84 divided by 4 = 21

OpenStudy (nincompoop):

how about 4t/4?

OpenStudy (anonymous):

1

OpenStudy (nincompoop):

no...

OpenStudy (nincompoop):

4/4 is 1 but you have t

OpenStudy (anonymous):

t = 21

OpenStudy (anonymous):

t > 21

OpenStudy (nincompoop):

NOT EQUAL. we are using inequality greater than or equal to, because in the original problem it says: spent no less than 84 dollars...

OpenStudy (anonymous):

\[t \ge 21\]

OpenStudy (nincompoop):

awesome

OpenStudy (anonymous):

[\[Tge21\]